If line k passes through the point (7,1) and is parallel to the graph of the line whose equation is y= 4x - 3, then what would the equation be?
Answer by Earlsdon(6294)

Since you don't specify the required form of the equation, let's use the point-slope form: (y - y1) = m(x - x1)
For this you will need a point on the line, given as (7, 1) and you will need the slope of the line. Note that parallel lines have identical slopes, so the slope of the line k will be the same as the slope of the line whose equation you are given. The given equation is in the slope-intercept form, y = mx + b so you can read the slope directly: m = 4.
So, the equation, in point-slope form, for line k would be: (y - 1) = 4(x - 7)